Titanium (IV) oxide is the main pigment in white and other light – colored paints. How many grams of titanium(IV) oxide (80) will produce from 18.0 g of oxygen (16) react with Ti to form titanium(IV) oxide?


1
Expert's answer
2020-12-17T07:34:32-0500

Since the molecular mass of oxygen is given as 16, It is assumed that oxygen is atomic. If the assumption is wrong, divide the final answer by 2. "Ti + 2O \rightarrow TiO_2"

"n= m\/M"

"M(O) = 16\frac{g}{mol}"

"M(TiO_2) = 80\frac{g}{mol}"

"n(O) = n(TiO_2) =\frac {18\ g }{16\frac{g}{mol}} = 1. 125 \ mol"

"m = n*M"

"m(TiO_2) = 1.125\ mol * 80\frac{g}{mol} =90\ g"

Answer: "m(TiO_2) = 90\ g"